# Spatialyze: A Geospatial Video Analytics System with Spatial-Aware Optimizations

### v3

@@ Abstract.
{NoHyper}\$*\$\$*\$footnotetext: Both authors contributed equally to this research.
@@ 1. Introduction
Geospatial videos are videos in which the location and time that they are shot are known. From surveillance footage, autonomous vehicle (AV) cameras to police body cameras, such videos are prevalent in our daily lives. While the volume of such data has grown tremendously (Wright, 2023; Sparks, 2023), we still lack end-to-end systems that can process and query such data effectively. The rise of machine learning (ML) in recent years has aggravated this issue, with the latest deep learning models capable of carrying out various computer vision tasks, such as object detection (Girshick et al., 2014; Jocher et al., 2022; Girshick, 2015; Redmon and Farhadi, 2018; Ren et al., 2016), multi-objects tracking (Bewley et al., 2016; Du et al., 2023; Wojke et al., 2017; Broström, 2022), image depth estimation (Godard et al., 2019), etc.
@@
All these trends have made geospatial video analytics computationally intensive. For instance, a \$2299 Nvidia T4 GPU takes 34 seconds on average to execute a simple workflow of object detection, tracking, and image depth estimation on a 20-second 12-fps video. Modern autonomous driving datasets, such as nuScenes (Caesar et al., 2019), contain 6000 such videos. Running the workflow mentioned above will take 3 full days to run on the entire dataset.
@@
To make matters worse, the lack of programming frameworks and data management systems for geospatial videos has made it challenging for end users to specify their workflows, let alone running them efficiently. Consider a data journalist writing an article on police misconduct and would like to examine the footage collected on police body cameras to look for specific behavior, e.g., police chasing after a suspect. Given today’s technology, they will either need to watch all the footage themself or string together various ML models for video analytics (Kang et al., 2017; Bastani et al., 2020; Bastani and Madden, 2022; Xu et al., 2022) and video processing libraries (e.g., OpenCV (Bradski, 2000), FFmpeg (Tomar, 2006)) to construct their workflow. Sadly, the former is simply infeasible given the amount of video data collected, while the latter requires deep programming expertise that most users do not possess.
@@
We see an opportunity to build a programming framework that bridges video processing and geospatial video analytics. Specifically, given that users constructing workflows on geospatial videos are typically interested in the time and location where such videos are taken, we should exploit the metadata stored in geospatial videos to optimize such workflows. For instance, a car is at (1810.39, 1221.71, 0.90) in the Boston Seaport map, which is at an intersection, at 2:33:59.112 PM of September 18, 2018. Each object inside a video captures a real-world object; therefore, this object also inherits physical behavior that its real-world counterpart would have had. For example, when users look for a car at a road intersection, we know that a road intersection need to be visible in the video frames that the car of interest is visible; therefore, video frames without a visible road intersection, determined with geospatial metadata, need not be processed with expensive ML functions. We leverage this already existing geospatial metadata, the inherited physical behaviors, and users’ queries to speed up video processing.
@@
To leverage this insight, we present Spatialyze, a system for geospatial video analytics. To make it easy for users to specify their geospatial video workflows, Spatialyze comes with a conceptual data model where users create and ingest videos into a “world,” and users interact with the world by specifying objects (e.g., cars) and scenarios (e.g., cars crossing an intersection) of interest via Spatialyze’s domain-specific language (DSL) embedded in Python called S-Flow. Spatialyze then efficiently executes the workflow by leveraging various spatial-aware optimization techniques that make use of existing geospatial metadata and assumptions based on inherited physical behavior of objects in the videos. For instance, Spatialyze’s Road Visibility Pruner uses road’s visibility as a proxy for objects’ visibility to prune out video frames. The Object Type Pruner then prunes out objects that are not of interest to the users’ workflow. Spatialyze’s 0pt speeds up object 3D location estimation by replacing a computationally expensive ML-based approach with a geometry-based approach. Finally, the Exit Frame Sampler prunes out unnecessary video frames based on the inherited physical behavior of vehicles and traffic rules. Spatialyze’s optimizations are all driven by various geospatial metadata and real-world physical behavior, with a goal to reduce the number of video frames to be processed and ML operations to be invoked, thus speeding up video processing runtime.
@@
As we are unaware of end-to-end geospatial video analytics systems, we evaluate different parts of Spatialyze against state-of-the-art (SOTA) video analytic tools (OTIF (Bastani and Madden, 2022), VIVA (Romero et al., 2022), EVA (Xu et al., 2022)), a geospatial data analytic tool (nuScenes devkit (Caesar et al., 2019)), and an aerial drone video sensing platform (SkyQuery (Bastani et al., 2021)) on different geospatial video workflows. We are up to 7.3$\times$ faster than EVA in geospatial object detection workflow, 1.06-2.28$\times$ faster than OTIF in object tracking, 1.68$\times$ faster than VIVA, 1.18$\times$ faster than SkyQuery in geospatial object tracking workflow, and 117-716$\times$ faster than nuScenes in geospatial data analytics. We also evaluate each of our optimization techniques by executing each of the queries with and without each of the techniques, shown in the results, achieving 2.5-5.3$\times$ speed up with 83.4-97.1% accuracy on 3D object tracking.
@@
In sum, we make following contributions. In § 4, we present a conceptual data model and a DSL (S-Flow) for geospatial video analytics. Using S-Flow, users construct their analytic workflows simply by following our build-filter-observe paradigm. In § 5, we describe the design of Spatialyze, a fast and extensible systems for processing arbitrary-length geospatial videos in S-Flow. In § 6, we develop 4 optimization techniques that leverage geospatial metadata embedded in videos and inherited physical behavior of objects to speed up geospatial video processing in Spatialyze’s geospatial video analytic workflow. In § 7, we implemented our techniques in Spatialyze and evaluated it against other SOTA video processing techniques as well as ablation studies, using real-world videos from nuScenes (Caesar et al., 2019), VIVA (Romero et al., 2022), and SkyQuery (Bastani et al., 2021).
@@ 2. Related Work
Recent use cases for querying geospatial video data (autonomous driving (Kim et al., 2021b), surveillance (Griffin and Miller, 2008), traffic analysis (Gloudemans and Work, 2021), and transshipment activities (McDonald et al., 2021; Park et al., 2020)) have driven attention to building geospatial DBMSs for data storage and retrieval. VisualWorldDB (Haynes et al., 2020) allows users to ingest video data from diverse sources and make them queryable as a single multidimensional object. It optimizes query execution by jointly compressing overlapping videos and reusing results from them. Apperception (Ge et al., 2021) provides a Python API for users to input geospatial videos, organizeing them into a four-dimensional data model, and query video data. While Apperception focuses on organizing and retrieving geospatial video data, it falls short in optimizing query execution. In contrast, Spatialyze provides a comprehensive geospatial video analytic workflow interface for declaratively specifying videos of interest. It further introduces new query optimization techniques by leveraging the inherent geospatial properties of videos.
@@
Researchers have proposed techniques to efficiently detect, track, and estimate 3D locations of objects from videos. Monodepth2 (Godard et al., 2019) estimates per-pixel image depth data using only monocular videos. EVA (Xu et al., 2022) identifies, materializes, and reuses expensive user-defined functions (UDF) on videos for exploratory video analytics with multiple refined and reused queries. However, EVA’s optimization techniques are aimed at scenarios where multiple queries use a high degree of overlapping UDFs. OTIF (Bastani and Madden, 2022) is a general-purpose object tracking algorithm. It uses a segmentation proxy model to determine if frames or regions of frames contain objects that need to be processed with an object detector, along with a recurrent reduced-rate tracking method that speed up tracking by reducing frame rate. VIVA (Kang et al., 2022; Romero et al., 2022) is a video analytic system that allows users to specify relationship opportunities of replacing or filtering one ML model with another. It uses these relationships to speed-up video processing time. However, it does not optimize video processing with geospatial metadata by default, which leaves substantial performance gains unexploited, as our experiments show. Complexer-YOLO (Simon et al., 2019) and EagerMOT (Kim et al., 2021a) propose one-stop machine learning models for efficiently tracking 3D objects using LiDAR data, which is not always available compared to cameras and GPS sensors. In addition, 3D sensing data is large. For instance, storing 850 20-second-long scenes of LiDAR sensor data in nuScenes requires 162 GB, while videos of the same scenes only take up 92 GB. Spatialyze overcomes this limitation by not requiring such information to process geospatial workflows, and instead readily available geospatial metadata to optimize user workflows. In addition, one of our optimization techniques estimates objects’ 3D location 192$\times$ faster than the approach that uses Monodepth2 on average. SkyQuery (Bastani et al., 2021) is a drone video sensing platform. It also provides a tool for querying and visualizing sensed aerial drone videos, but does not optimize video query processing.
@@ 3. Usage Scenario
In this section, we give an overview of Spatialyze using an example. Suppose a data journalist is interested in investigating AV footage reports, with the goal to find video snippets where other vehicles crash into the drivers’ vehicle at a road intersection.
@@
Naively, the data journalist would watch all the videos individually to identify parts of videos that are relevant. Even if the journalist has programming skills, they still need to manually invoke ML models to process large volumes of geospatial videos, and possibly optimize model inference by reducing the number of frames to be processed. They need to then merge these results with geospatial metadata (e.g., road information) to extract vehicle geospatial data, and write queries to isolate objects of interest and link them to the geospatial objects for a cohesive output. This process demands high coding expertise, is slow, tedious, and error-prone.
@@
Spatialyze streamlines this process. Instead of manually writing code to run ML models and extract geospatial information, journalists can use Spatialyze to easily construct such workflows. With Spatialyze, they can construct an analytic workflow with using S-Flow’s 3-step paradigm as shown in Fig. 1 to: Build a world, by integrating the video data with the geospatial metadata (e.g., road network and camera configurations); Filter the video for parts of interest, by describing relationships between objects (humans, cars, trucks) and their surrounding geospatial environment (lane, intersection); Observe the filtered video parts, in this case, by saving all the filtered video parts into video files for further examination.
@@
Build a world. In line LABEL:line:world, the user creates a world which represents a geospatial virtual environment to be discussed in § 4. Line LABEL:line:road then initializes the world by loading roadNetwork data, in this case, a set of the Boston Seaport’s road segments. Lines LABEL:line:gsvideo-LABEL:line:end-build load videos with their camera configurations (locations, rotations, and intrinsics (Zhang, 2014)) and associate them to create geospatial videos added into the world. Finally, in line LABEL:line:end-build, they have constructed a data-rich world with video cameras, and road segments, all related spatially and temporally.
@@
Filter for the video parts of interest. From the data journalist’s perspective, they can interact with objects and their trackings from each video added to the world through object() with the world initialized. Here, our journalist wants to find video snippets where other vehicles crash into the driver’s vehicle at a road intersection. They do so by adding a filter to the world to only include objects that match the provided predicates: that they are cars or trucks (line LABEL:line:pred-type), are within 50 meters of the camera (line LABEL:line:pred-distance), are at an intersection (line LABEL:line:pred-contain), and are moving towards the camera (in line LABEL:line:pred-heading). After filter is executed at line LABEL:line:pred-heading, the world will contain only cars or trucks on intersections that are moving towards the camera, as desired.
@@
Observe the filtered video parts. After applying the filter, the world has fewer objects compared to when it was initialized. When the journalist calls saveVideos() in line LABEL:line:observe, Spatialyze saves only video snippets with objects in the current world, significantly shortening the total video length. This allows the journalist to focus only on relevant segments with highlighted objects of interest. As shown, S-Flow provides a declarative interface where users do not specify “how” to find the video snippets of interest; instead, they only need to describe “what” their video snippets of interest look like, and Spatialyze will optimize accordingly.
@@
We next discuss Spatialyze in detail (in § 5) and how it optimizes workflow execution (in § 6) by leveraging video metadata and the physical behavior of real-world objects.
[…]
world.geogConstruct(type=’intersection’) )
[…]
Figure 1. Geospatial Video Analytic Workflow with Spatialyze
@@ 4. Constructing Geospatial Video Analytic Workflows
Spatialyze comes with a DSL called S-Flow embedded in python for users to construct their geospatial video data workflow. In this section, we describe S-Flow with Spatialyze’s conceptual data model.
@@ 4.1.1. World
As discussed in § 3 and in line LABEL:line:world of Fig. 1, users construct a World as a first step of S-Flow. Inspired by VisualWorldDB, a World represents a geospatial virtual environment which encompasses Geographic Constructs and Movable Objects, all coexisting with spatial relationships. These relationships, combined with the S-Flow, enable users to describe the videos of interest effectively.
The World also reflects the physical world, where Movable Objects exhibit realistic behaviors, like cars following lanes and adhering to speed limits. We refer to these phenomena as inherited physical behaviors of Movable Objects. We will discuss how Spatialyze leverage these behaviors to optimize video processing in § 6.

### v4

@@ Abstract.
PVLDB Reference Format:
PVLDB, 17(9): 2136 - 2148, 2024.
doi:10.14778/3665844.3665846 This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 17, No. 9 ISSN 2150-8097.
doi:10.14778/3665844.3665846
PVLDB Artifact Availability:
The source code, data, and/or other artifacts have been made available at https://spatialyze.github.io.
@@ 1. Introduction
Geospatial videos record the locations and periods of the recordings in addition to the visual information. Such videos are prevalent in our daily lives, from surveillance footage to autonomous vehicle (AV) cameras to police body cameras. While the volume of such data has grown tremendously (Wright, 2023; Sparks, 2023), we still lack end-to-end systems that can process and query such data effectively. The rise of machine learning (ML) in recent years has aggravated this issue, with the latest deep learning models capable of carrying out various computer vision tasks, such as object detection (Girshick et al., 2014; Jocher et al., 2022; Girshick, 2015; Redmon and Farhadi, 2018; Ren et al., 2016), multi-objects tracking (Bewley et al., 2016; Du et al., 2023; Wojke et al., 2017; Broström, 2022), image depth estimation (Godard et al., 2019), etc.
@@
All these trends have made geospatial video analytics computationally intensive. For instance, a \$2,299 NVIDIA T4 GPU takes 34 seconds on average to execute a simple workflow of object detection, tracking, and image depth estimation on a 20-second 12-fps video. Modern autonomous driving datasets, such as nuScenes (Caesar et al., 2019), contain 6000 such videos. Running the workflow mentioned above will take 3 full days to run on the entire dataset.
@@
To make matters worse, the lack of programming frameworks and data management systems for geospatial videos has made it challenging for end users to specify their workflows, let alone run them efficiently. For instance, a data journalist writing an article on self-driving car failures would like to examine the footage collected on car cameras to look for a specific behavior, e.g., two cars crossing at an intersection. They will either need to watch all the footage themself or string together various ML models for video analytics (Kang et al., 2017; Bastani et al., 2020; Bastani and Madden, 2022; Xu et al., 2022) and video-processing libraries (e.g., OpenCV (Bradski, 2000), FFmpeg (Tomar, 2006)) to construct their workflow. Sadly, the former is infeasible given the amount of video data collected, while the latter requires programming expertise that most users do not possess.
@@
We aim to build a programming framework that bridges video processing and geospatial video analytics. Specifically, as users constructing workflows on geospatial videos are typically interested in the time and location where such videos are taken, and such information is often stored as video metadata, we exploit that to optimize such workflows. Objects captured in videos inherit the physical behavior of real-world objects. For example, when users look for a car at a road intersection, the intersection must be visible in the video frames that the car of interest is visible; therefore, video frames without a visible intersection, which can be determined using geospatial metadata, can be skipped during query processing. We leverage such existing geospatial metadata, the inherited physical behaviors, and users’ queries to speed up video processing.
@@
Leveraging this insight, we present Spatialyze, a system for geospatial video analytics. To make it easy for users to specify their geospatial video analytics workflows, Spatialyze comes with a conceptual data model where users create and ingest videos into a “world,” and users interact with the world by specifying objects (e.g., cars) and scenarios (e.g., cars at an intersection) of interest via Spatialyze’s S-Flow, a domain-specific language (DSL) embedded in Python. Spatialyze then efficiently executes the workflow by leveraging various spatial-aware optimization techniques that make use of existing geospatial metadata and assumptions based on the inherited physical behavior of objects in the videos. For instance, Spatialyze’s Road Visibility Pruner uses the road’s visibility as a proxy for objects’ visibility to prune out video frames. The Object Type Pruner then prunes out objects that are not of interest to the users’ workflow. Spatialyze’s 0pt speeds up object 3D location estimation by replacing a computationally expensive ML-based approach with a geometry-based approach. Finally, the Exit Frame Sampler prunes out unnecessary video frames based on the inherited physical behavior of vehicles and traffic rules. Spatialyze’s optimizations are all driven by various geospatial metadata and real-world physical behavior, with the goal of reducing the number of video frames to be processed and ML operations to be invoked, thus speeding up the video processing runtime.
@@
As we are unaware of end-to-end geospatial video analytics systems, we evaluate different parts of Spatialyze against state-of-the-art (SOTA) video analytic tools (OTIF (Bastani and Madden, 2022), VIVA (Romero et al., 2022), EVA (Xu et al., 2022)), a geospatial data analytic tool (nuScenes devkit (Caesar et al., 2019)), and an aerial drone video sensing platform (SkyQuery (Bastani et al., 2021)) on different geospatial video analytics workflows. We are up to 7.3$\times$ faster than EVA in geospatial object detection workflow, 1.06-2.28$\times$ faster than OTIF in object tracking, 1.68$\times$ faster than VIVA, 1.18$\times$ faster than SkyQuery in geospatial object tracking workflow, and 117-716$\times$ faster than nuScenes in geospatial data analytics. We also evaluate each of our optimization techniques by executing each of the queries with and without each of the techniques, shown in the results, achieving 2.5-5.3$\times$ speed up with 83.4-97.1% accuracy on 3D object tracking.
@@
In sum, we make the following contributions. In § 4, we present a conceptual data model and a DSL (S-Flow) for geospatial video analytics. Using S-Flow, users construct their analytic workflows simply by following our build-filter-observe paradigm. In § 5, we describe the design of Spatialyze, a fast and extensible system to process geospatial videos of arbitrary length in S-Flow. In § 6, we develop 4 optimization techniques that leverage geospatial metadata embedded in videos and inherited physical behavior of objects to speed up geospatial video processing in Spatialyze’s geospatial video analytic workflow. In § 7, we implemented our techniques in Spatialyze and evaluated it against other SOTA video processing techniques, as well as ablation studies, using real-world videos from nuScenes (Caesar et al., 2019), VIVA (Romero et al., 2022), and SkyQuery (Bastani et al., 2021).
@@ 2. Related Work
Recent use cases for querying geospatial video data (autonomous driving (Kim et al., 2021b), surveillance (Griffin and Miller, 2008), traffic analysis (Gloudemans and Work, 2021), and transshipment activities (McDonald et al., 2021; Park et al., 2020)) have led to the development of geospatial DBMSs for data storage and retrieval. VisualWorldDB (Haynes et al., 2020) allows users to ingest video data from various sources and make them queryable as a single multidimensional object. It optimizes query execution by jointly compressing overlapping videos and reusing results from them. Apperception (Ge et al., 2021) provides a Python API for users to input geospatial videos, organize them into a four-dimensional data model, and query video data. While Apperception focuses on organizing and retrieving geospatial video data, it falls short of optimizing query execution. In contrast, Spatialyze provides a comprehensive geospatial video analytic workflow interface for declaratively specifying videos of interest. In addition, it introduces new query optimization techniques by leveraging the inherent geospatial properties of videos.
@@
Researchers have proposed techniques to efficiently detect, track, and estimate 3D locations of objects from videos. Monodepth2 (Godard et al., 2019) estimates per-pixel image depth data using only monocular videos. EVA (Xu et al., 2022) identifies, materializes, and reuses expensive user-defined functions (UDF) on videos for exploratory video analytics with multiple refined and reused queries. However, EVA’s optimization techniques aim at scenarios where multiple queries repeatedly invoke overlapping UDFs. OTIF (Bastani and Madden, 2022) is a general-purpose object tracking algorithm. It uses a segmentation proxy model to determine if frames or regions of frames contain objects that need to be processed with an object detector, along with a recurrent reduced-rate tracking method that speeds up tracking by reducing frame rate. VIVA (Kang et al., 2022; Romero et al., 2022) is a video analytic system that allows users to specify relationship opportunities of replacing or filtering one ML model with another. It uses these relationships to speed up the video processing time. However, it does not optimize video processing with geospatial metadata by default, leaving substantial performance gains unexploited, as our experiments show. Complexer-YOLO (Simon et al., 2019) and EagerMOT (Kim et al., 2021a) propose one-step ML models for efficiently tracking 3D objects using LiDAR data, which is larger and not always available compared to cameras and GPS sensors information. Spatialyze overcomes this limitation by not requiring such information to process geospatial workflows, but instead using readily available geospatial metadata to optimize user workflows. In addition, one of our optimization techniques estimates objects’ 3D location 192$\times$ faster than the approach that uses Monodepth2 on average. SkyQuery (Bastani et al., 2021) is a drone video sensing platform. It also provides a tool for querying and visualizing sensed aerial drone videos but does not optimize video query processing.
@@ 3. Usage Scenario
In this section, we give an overview of Spatialyze using an example. Suppose a data journalist is interested in investigating AV footage reports, and the goal is to find video snippets of other vehicles that crash into the driver’s vehicle at a road intersection.
@@
Naively, the data journalist would watch all the videos individually to identify parts of the videos that are relevant. Even with programming skills, they still need to manually invoke ML models to process large volumes of geospatial videos, and possibly optimize model inference by reducing the number of frames to be processed. They need to then merge these results with geospatial metadata (e.g., road information) to extract vehicle geospatial data, write queries to isolate objects of interest, and link them to the geospatial objects for a cohesive output. This process demands high coding expertise and is slow, tedious, and error-prone.
@@
Spatialyze streamlines this process. Instead of manually writing code to run ML models and extract geospatial information, journalists can use Spatialyze to construct such workflows easily. With Spatialyze, they can construct an analytic workflow with S-Flow’s 3-step paradigm as shown in § 3 to: Build a world, by integrating the video data with the geospatial metadata (e.g., road network and camera configurations); Filter the video for parts of interest, by describing relationships between objects (humans, cars, trucks) and their surrounding geospatial environment (lane, intersection); Observe the filtered video parts, in this case, by saving all the filtered video parts into video files for further examination.
@@
Build a world. In line 3, the user creates a world which represents a geospatial virtual environment to be discussed in § 4. Line 3 then initializes the world by loading roadNetwork data, in this case, a set of the Boston Seaport’s road segments. Lines 3-3 load videos with their camera configurations (locations, rotations, and intrinsics (Zhang, 2014)) and associate them to create geospatial videos added into the world. Finally, in line 3, they have constructed a data-rich world with video cameras, and road segments, all related spatially and temporally.
@@
Filter for the video parts of interest. From the data journalist’s perspective, they can interact with objects and their trackings from each video added to the world through object() with the world initialized. Here, our journalist wants to find video snippets where other vehicles crash into the driver’s vehicle at a road intersection. They do so by adding a filter to the world to only include objects that match the provided predicates: that they are cars or trucks (line 3), are within 50 meters of the camera (line 3), are at an intersection (line 3), and are moving towards the camera (in line 3). After filter is executed at line 3, the world will contain only cars or trucks on intersections that are moving towards the camera, as desired.
@@
Observe the filtered video parts. After applying the filter, the world has fewer objects than when it was initialized. When the journalist calls saveVideos() in line 3, Spatialyze saves only video snippets with objects in the current world, significantly shortening the total video length. This allows the journalist to focus only on relevant segments with highlighted objects of interest. As shown, S-Flow provides a declarative interface where users do not specify “how” to find the video snippets of interest; instead, they only need to describe “what” their video snippets of interest look like, and Spatialyze will optimize accordingly.
@@
We next discuss Spatialyze in detail (in § 5) and how it optimizes workflow execution (in § 6) by leveraging videos’ geospatial metadata and the physical behavior of real-world objects. {listing}[!ht]{minted}[ linenos, mathescape, escapeinside=!!, fontsize=]python world = World() # OR World(detect=CstmDetector, track=CstmTracker) world.addGeogConstructs(RoadNetwork(’road-network/boston-seaport/’)) !! world.addVideo(GeospatialVideo(Video(’v0.mp4’), Camera(’c0.json’))) !! world.addVideo(GeospatialVideo(Video(’v1.mp4’), Camera(’c1.json’))) !! obj, cam, intersection = ( world.object(), world.camera(), !! world.geogConstruct(type=’intersection’) ) world.filter( ((obj.type == ’car’) — (obj.type == ’truck’)) !! !! & (distance(obj, cam) ¡ 50) contains(intersection, obj) !! !! headingDiff(obj, cam, between=[135, 225]) ) !! world.saveVideos(’output_videos/’, addBoundingBoxes=True) !! Geospatial Video Analytic Workflow with Spatialyze
@@ 4. Constructing Geospatial Video Analytic Workflows
Spatialyze comes with S-Flow, a DSL embedded in Python for users to construct their geospatial video data workflow. In this section, we describe S-Flow with Spatialyze’s conceptual data model.
@@ 4.1.1. World
As discussed in § 3 and in line 3 of § 3, users construct a World as a first step of S-Flow. Inspired by VisualWorldDB, a World represents a geospatial virtual environment which encompasses Geographic Constructs and Movable Objects, all coexisting with spatial relationships. These relationships, combined with the S-Flow, enable users to describe the videos of interest effectively.
The World also reflects the physical world, where Movable Objects exhibit realistic behaviors, like cars following lanes and adhering to speed limits. We refer to these phenomena as inherited physical behaviors of Movable Objects. We will discuss how Spatialyze leverages these behaviors to optimize video processing in § 6.